Determine the value of sin2x given that tan^2x=81/49 and Pi<x<3pi/2


1
Expert's answer
2021-01-21T19:23:08-0500

"tan²x = \frac{81}{49}"

=> "tan x = \frac{9}{7}"

"x = tan^{-1} (\frac{9}{7})
x = 52.1250^{0}"


Since tan is positive in both 1st and 3rd quadrant, that implies


In 1st quadrant

x = 52.1250°


In the 3rd quadrant


x = 180 + 52.1250° = 232.1250°


But "π<x<\frac{3π}{2}"


Therefore, x = 232.1250°


"Sin 2x = Sin 2(232.1250) = 0.9692"